OpenStudy (anonymous):

Calculate the work needed for a 65 kg person to climb through 4.0 m on the surface of (a) the Earth and (b) the Moon

OpenStudy (anonymous):

Work = Force * Displacement Force= Mass * acceleration due to gravity Given : Mass of a Person = 65 Kg accelaration due to gravity on the surface of the earth = 9.8m/s2 Displacement of the body = 4.0 m Work done = 65 Kg * 9.8 m/s2 * 4.0 m = 2548 Joules When the body is taken to the surface of the moon----- accleration due to gravity on the surface of themoon = 1.60 m/s2 Work done = 65 Kg * 1.60 m/s2 * 4.0m = 416 Joules Source: http://www.chegg.com/homework-help/questions-and-answers/calculate-work-needed-65-kg-person-climb-40-mon-surface-earth-b-moon-g-160-ms-1--q318756

OpenStudy (anonymous):

Work = change in potential energy Potential energy = mgh.